A hole of cross-sectional area 100 c m 2 is cut in a piece of steel at 20 o C . What is the area of the hole if the steel is heated from 20 o C to 100 o C ? The coefficient of linear expansion of steel is 1.10 × 10 − 5 o C


1
Expert's answer
2021-11-08T17:03:05-0500

We can find the area of the hole if the steel is heated from "20^{\circ}C" to "100^{\circ}C" as follows:


"\Delta A=2\alpha A_0\Delta T,""A-A_0=2\alpha A_0\Delta T,""A=A_0(1+2\alpha\Delta T),""A=100\ cm^2\times(1+2\times1.10\times10^{-5}\ (^{\circ}C)^{-1}\times80^{\circ}C),""A=100.2\ cm^2."